# A Unified Approach to Extract Intepretable Rules from Tree Ensembles via Integer Programming

### v1

@@ 3.2 Decision trees
We consider standard supervised classification and regression problems, with a given a training dataset of i.i.d. samples , where data points are expressed as vectors containing tabular features or time series of length . In the case of classification task, targets represent a categorical variable, where is a set of class labels. In the case of regression task, targets represent a numerical variable. The given dataset is used to construct a predictor that maps each data point to the codomain of the corresponding task. A binary decision tree of maximal depth is a predictor that recursively divides the feature space into at most non-overlapping partitions called leaf nodes. Each partition is defined by an ordered collection of disjunctive splittings called branch nodes. Splittings are defined by a tuple , for each , where is the set of branch nodes. The first element of the tuple represents the feature selector, whereas the second element corresponds to the threshold value. Each splitting compares the selected feature to the threshold value. If a data point exhibits a value lower or equal than the threshold, it is assigned to the left child of the branch node; otherwise, it is assigned to the right one. In the case of tabular data, the feature selector is the scalar product between point and one canonical base of the same feature space . Figure 1 depicts an example of a binary decision tree of maximal depth for the classification of tabular data. It contains branch nodes and leaf nodes . The root node selects the 10th feature of each point contained in the dataset, and compares each value to the threshold . The other branch nodes operate in a similar way. Each leaf node is associated to a colour representing one of two class labels in .
Shapelets-based tree.
@@
If points contained in the dataset are time series, the first element of each splitting represents an extractor of temporal features. Without loss of generality, we consider as temporal feature the distance (1) computed between time series and meaningful subsequences sampled from the dataset. We show how to use data-mined shapelets as subsequences to learn a decision tree for time series classification [66, 10]. As an example, we use the ItalyPowerDemand dataset from the UCR repository [24]. The training set includes 67 time series, each 24 units long, representing twelve monthly electrical power demand patterns from Italy. The classification task is to distinguish days between October and March from those between April and September, which correspond to the two class labels in . The top two subplots in Figure 2 show the training data for the two classes. Using one of the available data mining approaches [57, 32], it is possible to extract shapelets and from the training data. These shapelets are depicted in dark green in the rightmost part of Figure 2. By computing distance (1) between each time series and each of the two shapelets and , we can represent each time series as a point in the scatter plot in Figure 2. The horizontal axis represents dist, while the vertical axis represents dist. Finally, using the two distances as features for each time series, we can build a decision tree that partitions the feature space associating threshold values and with shapelets and , respectively. Figure 3 shows a shapelets-based decision tree of maximal depth , with branch nodes and leaf nodes . Threshold values correspond to those shown in the scatter plot in Figure 2.
Rule extraction.
@@
We define as rule the splitting of a given branch node paired with the left () or right () sign. For instance, we indicate with the rule that a given point has to satisfy to get to the left child of the root node, that is whether or not the inequality holds. Starting from the root node, each point is recursively evaluated at branch nodes until it is finally assigned to a leaf node to get a prediction. In the case of a classification tree, the prediction corresponds to the most frequent class contained in the leaf node. For regression, the prediction is the average value of the target variables of the assigned data points. Since the path from the root node to each leaf node is unique, we can represent each leaf node as the ordered list of rules along its path
@@
where stands for the sign of each rule, while represents the parent node of leaf node . Since splittings are disjunctive, each point that satisfies all the rules of a list is uniquely assigned to the corresponding leaf node. Given a list of rules , we define the set that contains the corresponding splitting of each rule as follows
@@
As example, let us consider the decision tree depicted in Figure 1. The rule list corresponding to the first leaf node is
@@
Finally, we can represent a decision tree by all the paths it contains, expressing it as the set containing the corresponding rule lists
For instance, the decision tree in Figure 1 can be represented as the following set of rule lists
whereas the decision tree for time series data in Figure 3 can be represented as
@@ 3.3 Tree ensembles
A tree ensemble classification or regression method consists of learning multiple decision trees by bagging [13] or boosting [29], leveraging their aggregate predictions to obtain more accurate and robust results. The former approach consists of training a collection of decision trees with different samples bootstrapped from the same training dataset, obtaining a random forest [14]. Instead, in the latter approach trees are sequentially trained by exploiting the prediction error of the previous ones, obtaining boosted learners like XGBoost [20]. For sake of simplicity, in the reminder of this paper we focus on random forests as representative of tree ensemble methods. Exploiting the introduced formalism, we can express a random forest containing decision trees , with as their disjoint union, that is
In the classification setting, the prediction of a random forest corresponds to the majority vote of its decision trees. For regression tasks, the prediction is the average of all the predictions obtained by its decision trees. A generalized Random Shapelet Forest [38] is obtained by bagging decision trees whose feature selectors correspond to (1) applied to sampled shapelets. Similarly, a Time Series Forest [26] can be obtained by applying the same procedure choosing as feature one among mean, standard deviation and slope computed on sampled intervals. Thanks to the formalism we introduced, we can represent any random forest as a set of data-agnostic rule lists. Hereafter, we omit the arguments of the aforementioned mathematical objects to simplify the notation.
4 Our approach
In this section, we present and discuss the proposed method. Given a trained random forest containing a total of rule lists, our approach consists of selecting the optimal rule lists by solving a set partitioning problem. In this way, we can guarantee that each point is assigned to only one of the rule lists it satisfies, mimicking a decision tree and improving interpretability. The key idea is to select the best subset of rules according to their robustness to data perturbation and predictive power, measured by computing a rule-specific regularization and loss function, respectively.
@@ 4.1 Preprocessing step
Since set partitioning problems are NP-hard [40], we alleviate the computational burden of our approach by separating the computation of regularization and loss functions from the optimization problem. Thus, we characterize each rule list by three quantities that are computed during the following preprocessing step. We define the stability of the -th rule list as the sum of the Sørensen-Dice index [18, 41] computed on all the lists contained in , ignoring their inequality signs
that is the total amount of shared splittings between rule list and all the other lists in the forest. In other words, we measure how much a rule is replicated through the whole forest. A list made of rules that are consistently replicated inside the random forest is more stable than a list containing less frequent rules. By maximizing their stability, selected lists are less sensitive to data perturbation and they can more likely resemble a decision tree structure. This approach combines the effects of the fusion penalty of FIRE [42] and the occurrence frequency of SIRUS [6]. Indeed, the former penalty term promotes the learning of rules with more shared splittings, improving their interpretability. Instead, the latter method prunes the rules that entirely occur less than the frequency threshold parameter . In order to quantify the predictive power of the -th rule list, we compute its loss depending on the learning task as follows
where is the number of data points that satisfy the rule list, and is the number of data points of class among them. Furthermore, MSE corresponds to the Mean Squared Error (MSE) computed on the response values of data points that satisfy rule list . In other words, we quantify the loss of the -th rule list as the number of misclassified points in the case of classification tasks. Instead, we consider the MSE computed on assigned points as the loss of the -th rule list in the case of regression tasks. By computing loss values during this preprocessing step, we manage to easily handle non-linear and multiple regularization and loss functions without solving any optimization problem or tuning any parameter. After their computation, vectors and are normalized to the same scale. Lastly, we keep track of the assignment for each of the data point to each of the lists through the matrix , whose entries are computed as follows
@@ 5 Experiments and results
In this section, we present and discuss the experimental evaluation conducted to assess the performance of the proposed method against the state-of-the-art, for regression and classification tasks of tabular data and time series classification. To ensure a fair comparison, we selected as competing approaches the state-of-the-art methods that are clearly reproducible. We implemented all methods in Python 3.9, except for SIRUS, which was implemented in R 4.3.1. We ran all the experiments on the same machine equipped with a 3.3 GHz 8-core processor and 16 GB of memory. The proposed IP problems were solved through Gurobi 11.0 [34]. In order to compare multiple classifiers over the selected datasets, we perform the non-parametric Friedman test [25] combined with the signed-rank Wilcoxon-Holm post-hoc test [36], with a level of statistical significance equal to . According to Benavoli et al. [8], the Wilcoxon signed-rank test is more appropriate than the more common post-hoc Nemenyi test [49]. Details about benchmark datasets are contained in Table 2, Table 4 and Table 6. During data preprocessing, we applied one-hot encoding to tabular data with categorical features and discarded missing values. For regression tasks, we standardized the response vector. No preprocessing was applied to time series datasets.
@@ 5.1 Tabular regression
In this experiment, we compare the proposed method to the state-of-the-art about rule extraction for regression tasks of tabular data. In particular, competing methods are RuleFit [30], SIRUS [6] and NodeHarvest (NH) [46]. The experimental plan involves running each method on a selected benchmark of 22 datasets from the OpenML public repository [60]. Following the experimental plan of earlier works [42, 46], we extracted rules from a regression forest of 500 trees of depth 3 implemented through the scikit-learn package [48], and fixed the maximum amount of selected rule lists to 15. Since this quantity coincides with parameter , the proposed training procedure is not needed, and our method runs without validation or tuning steps. For competing methods, we respected the parameter tuning procedures described by their authors. We divided each dataset into 75/25 train-test split. We ran each regression task with 30 different random seeds, and we report average and standard deviation of the out-of-sample Mean Squared Error (MSE) in Table 2. For sake of clarity, we also report the results of the random forest (Full model). From obtained results, all methods achieves statistically equivalent performance (Friedman’s test -value ). Interestingly, all rule extraction methods manage to outperform the Full model, obtaining improved interpretability and performance. The median value of computing times of the proposed method is minutes, and the average runtime of the Gurobi IP solver is of the total computing time. Competing methods are more efficient than our approach. Nevertheless, under this setting, the latter operates as a parameter-free method.
@@
| Dataset | Our method | RuleFit | SIRUS | NH | Full model |
| Mercedes_Benz | 0.45 0.10 | 0.45 0.11 | 0.51 0.10 | 0.45 0.10 | 0.61 0.05 |
| Moneyball | 0.25 0.14 | 0.21 0.08 | 0.20 0.07 | 0.26 0.08 | 0.31 0.06 |
| abalone | 0.63 0.06 | 0.60 0.06 | 0.68 0.06 | 0.60 0.06 | 0.59 0.04 |
| autoMpg | 0.31 0.12 | 0.31 0.08 | 0.28 0.05 | 0.32 0.08 | 0.25 0.06 |
| bank32nh | 0.65 0.03 | 0.64 0.04 | 0.61 0.04 | 0.65 0.04 | 0.75 0.04 |
| boston | 0.34 0.11 | 0.28 0.06 | 0.29 0.09 | 0.31 0.06 | 0.31 0.03 |
| cpu_small | 0.09 0.01 | 0.08 0.01 | 0.11 0.01 | 0.08 0.01 | 0.12 0.01 |
| elevators | 0.52 0.03 | 0.54 0.05 | 0.64 0.04 | 0.53 0.03 | 0.53 0.06 |
| house_16H | 0.68 0.05 | 0.69 0.06 | 0.66 0.06 | 0.69 0.05 | 0.68 0.02 |
| houses | 0.51 0.01 | 0.52 0.03 | 0.56 0.02 | 0.51 0.01 | 0.53 0.11 |
| kin8nm | 0.41 0.02 | 0.68 0.02 | 0.65 0.03 | 0.68 0.05 | 0.66 0.06 |
| mtp | 0.77 0.03 | 0.73 0.04 | 0.74 0.03 | 0.75 0.06 | 0.73 0.03 |
| no2 | 0.58 0.13 | 0.58 0.12 | 0.54 0.10 | 0.60 0.13 | 0.59 0.10 |
| pol | 0.32 0.01 | 0.29 0.05 | 0.28 0.02 | 0.30 0.01 | 0.61 0.02 |
| puma32H | 0.42 0.01 | 0.45 0.08 | 0.56 0.06 | 0.42 0.01 | 0.86 0.02 |
| socmob | 0.45 0.25 | 0.31 0.11 | 0.37 0.12 | 0.37 0.25 | 0.49 0.21 |
| space_ga | 0.41 0.09 | 0.55 0.10 | 0.53 0.10 | 0.56 0.09 | 0.54 0.09 |
| stock | 0.09 0.01 | 0.11 0.03 | 0.24 0.08 | 0.10 0.01 | 0.11 0.02 |
| tecator | 0.05 0.11 | 0.05 0.02 | 0.08 0.02 | 0.04 0.02 | 0.20 0.01 |
| us_crime | 0.45 0.19 | 0.33 0.04 | 0.35 0.09 | 0.50 0.09 | 0.48 0.07 |
| wind | 0.38 0.01 | 0.37 0.03 | 0.39 0.02 | 0.36 0.03 | 0.33 0.01 |
| wine_quality | 0.69 0.04 | 0.75 0.04 | 0.78 0.04 | 0.75 0.02 | 0.75 0.04 |
| Average | 0.43 0.07 | 0.43 0.06 | 0.46 0.06 | 0.45 0.05 | 0.50 0.05 |

### v3

@@ 3.2 Decision trees
We consider standard supervised classification and regression problems, with a given a training dataset of i.i.d. samples , where data points are expressed as vectors containing tabular features or time series of length . In the case of classification task, targets represent a categorical variable, where is a set of class labels. In the case of regression task, targets represent a numerical variable. The given dataset is used to construct a predictor that maps each data point to the codomain of the corresponding task. A binary decision tree of maximal depth is a predictor that recursively divides the feature space into at most non-overlapping partitions called leaf nodes. Each partition is defined by an ordered collection of disjunctive splittings called branch nodes. Splittings are defined by a tuple , for each , where is the set of branch nodes. The first element of the tuple represents the feature selector, whereas the second element corresponds to the threshold value. Each splitting compares the selected feature to the threshold value. If a data point exhibits a value lower or equal than the threshold, it is assigned to the left child of the branch node; otherwise, it is assigned to the right one. In the case of univariate decision trees for tabular data, the feature selector is the scalar product between point and one canonical base of the same feature space . Figure 1 depicts an example of an univariate binary decision tree of maximal depth for the classification of tabular data. It contains branch nodes and leaf nodes . The root node selects the 10th feature of each point contained in the dataset, and compares each value to the threshold . The other branch nodes operate in a similar way. Each leaf node is associated to a colour representing one of two class labels in .
3.2.1 Shapelets-based trees
@@
If points contained in the dataset are time series, the first element of each splitting represents an extractor of temporal features. Without loss of generality, we consider as temporal feature the distance (1) computed between time series and meaningful subsequences sampled from the dataset. We show how to use data-mined shapelets as subsequences to learn a decision tree for time series interpretable classification [91, 13]. As an example, we use the ItalyPowerDemand dataset from the UCR repository [31]. The training set includes 67 time series, each 24 units long, representing twelve monthly electrical power demand patterns from Italy. The classification task is to distinguish days between October and March from those between April and September, which correspond to the two class labels in . The top two subplots in Figure 2 show the training data for the two classes. Using one of the available data mining approaches [78, 44], it is possible to extract shapelets and from the training data. These shapelets are depicted in dark green in the rightmost part of Figure 2. By computing distance (1) between each time series and each of the two shapelets and , we can represent each time series as a point in the scatter plot in Figure 2. The horizontal axis represents dist, while the vertical axis represents dist. Finally, using the two distances as features for each time series, we can build a decision tree that partitions the feature space associating threshold values and with shapelets and , respectively. Figure 3 shows a shapelets-based decision tree of maximal depth , with branch nodes and leaf nodes . Threshold values correspond to those shown in the scatter plot in Figure 2.
3.2.2 Rule extraction
@@
We define as condition the splitting of a given branch node paired with the left () or right () sign. For instance, we indicate with the condition that a given point has to satisfy to get to the left child of the root node, that is whether or not the inequality holds. Starting from the root node, each point is recursively evaluated at branch nodes until it is finally assigned to a leaf node to get a prediction. In the case of a classification tree, the prediction corresponds to the most frequent class contained in the leaf node. For regression, the prediction is the average value of the target variables of the assigned data points. Since the path from the root node to each leaf node is unique, we can represent each leaf node as a rule, that is the ordered collection of conditions along its path
@@
where stands for the sign of each condition, while represents the parent node of leaf node . Since splittings are disjunctive, each point that satisfies all the conditions of a rule is uniquely assigned to the corresponding leaf node. Given a rule , we define the set that contains all the corresponding splitting as follows
@@
As example, let us consider the decision tree depicted in Figure 1. The rule corresponding to the first leaf node is
@@
Finally, we can represent a decision tree by all the paths it contains, expressing it as the set containing the corresponding rules
@@ 3.3 Tree ensembles
A tree ensemble classification or regression model consists of learning multiple decision trees by bagging [16] or boosting [39], leveraging their aggregate predictions to obtain more accurate and robust results. The former approach consists of training a collection of decision trees with different samples bootstrapped from the same training dataset, obtaining a random forest [17]. Instead, in the latter approach trees are sequentially trained by exploiting the prediction error of the previous ones, obtaining boosted learners like XGBoost [25]. For sake of simplicity, in the reminder of this paper we focus on random forests as representative of tree ensemble models. Exploiting the introduced formalism, we can express a random forest containing decision trees , with as their disjoint union, that is
In the classification setting, the prediction of a random forest corresponds to the majority vote of its decision trees. For regression tasks, the prediction is the average of all the predictions obtained by its decision trees. A generalized Random Shapelet Forest [52] is obtained by bagging decision trees whose feature selectors correspond to (1) applied to sampled shapelets. Similarly, a Time Series Forest [33] can be obtained by applying the same procedure choosing as feature one among mean, standard deviation and slope computed on sampled intervals. With this formalism, we can represent any random forest as a set of rules.
3.4 Rule fidelity
Our goal is to extract a surrogate model from a trained random forest in the form of a set (or list) containing rules. By doing so, we aim to effectively retain the predictive power of the tree ensemble and faithfully represent its inner structure in an interpretable way. To measure the internal fidelity of a set of rules towards the tree ensemble, we introduce two novel measures of representativeness. The first one is the percentage of trees whose paths are represented by a list of rules. A tree is path-represented by a list of rules if . In other words, a tree of the random forest is path-represented by a list if the latter contains at least one of the paths from the root node of the tree to its leaf nodes. The second fidelity measure corresponds to the percentage of trees whose conditions (or branch nodes) are represented by a list of rules. A tree is node-represented by a list of rules if the latter contains at least one logical condition (or branch node) of the tree. In this way, we can measure how closely a list of rules resembles the random forest, depending on the amount of path-represented and node-represented trees. For example, the tree in Figure 1
is path-represented and node-represented by the rule
while it is only node-represented by the rule
as they only share the condition .
4 Methodology
In this section, we present and discuss the proposed method. Given a trained random forest containing a total of rules, our approach consists of selecting an optimal list of rules by solving a set partitioning problem. In this way, we can guarantee that each point is assigned to only one of the rules it satisfies, mimicking the partition of the feature space provided by decision trees and enhancing overall interpretability. The key idea is to select the best subset of rules according to their robustness to data perturbation and predictive power, measured by computing rule-specific regularization and loss function, respectively.
@@ 4.1 Preprocessing step
Since the set partitioning problem is NP-hard [58], we alleviate the computational burden of our approach by separating the computation of regularization and loss functions from the optimization problem. Therefore, we characterize each rule by three quantities that are computed during the following preprocessing step. We define the stability of the -th rule as the weighted sum of the Sørensen-Dice index [23, 59] computed over all the rules contained in the random forest, ignoring their inequality sign. Given the set of splittings in the form (3), for each , where is the amount of rules in the random forest, we compute the stability of rule as follows
where is the weight of the -th rule in case of weighted tree ensembles ( if unweighted). This measures the total amount of shared splittings between rule and all the other rules in the tree ensemble. As example, consider the following rules extracted from a random forest
with their corresponding sets of splittings
The stability of each rule is computed as follows
In other words, we measure how much a logical condition is replicated through the whole ensemble. A rule made of conditions that are consistently replicated inside the random forest is more stable than a rule containing less frequent conditions. By maximizing their stability, selected rules are less sensitive to data perturbation, and they can more likely resemble a decision tree structure. This approach combines the effects of the fusion penalty of FIRE [60] and the occurrence frequency of SIRUS [8]. Indeed, the former penalty term promotes the learning of rules with more shared splittings, improving their interpretability. Instead, the latter method prunes the rules that entirely occur less than the frequency threshold parameter . In order to quantify the predictive power of the -th rule, we compute its loss depending on the learning task as follows
where is the number of data points that satisfy the rule, is the number of data points of class among them, and MSE corresponds to the Mean Squared Error (MSE) computed on the response values of data points that satisfy rule . In other words, we quantify the loss of the -th rule as the number of misclassified points in the case of classification tasks. Instead, we consider the MSE computed on assigned points as the loss of the -th rule in the case of regression tasks. By computing loss and regularization values during the preprocessing step, we manage to easily handle non-linear and multiple functions. After their computation, vectors and are normalized to the same scale. Lastly, we keep track of the assignment for each of the data point to each of the rules through the matrix , whose entries are computed as follows
@@ 4.3 Validation procedure
that is the number of rules (or leaf nodes) contained in the smallest tree of the ensemble. This approach yields lower bounds that are close to the exact values, as shown in Figure 4. To compute a heuristic value for the upper bound , we can construct a decision tree with low cost complexity on the input data and count the number of generated leaf nodes. We note that obtained results can significantly vary depending on the dataset. After computing bounds, we solve Problem (4.2) for each value of over the training set. We remark that reoptimizing the same IP problem with a different right-hand side is computationally efficient, since modern solvers can exploit effective warm start techniques [88]. Finally, we select the value of that achieves the lowest loss on the validation set.
4.4 Inference procedure
One of the major differences between the proposed method and existing works on rule extraction consists of the lack of weighted rules. Indeed, our approach generates a list of rules that corresponds to a set partition, without incorporating any weighting strategy. We argue that this approach is more interpretable than competing ones, as it resembles a decision tree and it partitions the training data into a rule-based clustering structure, which is considered a challenging problem in the field of interpretable machine learning [20]. However, the application of an unweighted list of rules may fail on unseen data points, especially if they significantly differ from the training data. This can result in out-of-sample data falling into multiple rules or none at all, preventing the proposed method from generating a meaningful prediction. We remark that, in rule extraction for black-box models, the purpose is to explain the predictions obtained from a black-box model [94]. Therefore, if the surrogate model fails to obtain an individual prediction as in the case of multiple or absent rules, the unseen data point can still be predicted by the tree ensemble [50]. Nonetheless, we propose a simple heuristic to address the issue. If an unseen data point falls within two or more rules, we apply the rule that covers the largest amount of training data. Instead, if an unseen data point does not fall within any rules of the list, the prediction is the average of the response vector in case of regression, or the majority class in case of classification.
@@ 5 Experiments and results
In this section, we present and discuss the experimental evaluation conducted to assess the performance of the proposed method against the state-of-the-art, for regression and classification tasks of tabular data and time series classification. To ensure a fair comparison, we selected as competing approaches the state-of-the-art methods that are reproducible. For the same reason, we excluded FIRE from the comparison, since it does not provide a fixed upper limit on the number of extracted rules. We discuss more in detail in A the challenges of establishing a fair comparison for rule extraction methods. We implemented all methods in Python 3.9, except for SIRUS, which was implemented in R 4.3.1. We ran all the experiments on the same machine equipped with a 3.3 GHz 8-core processor and 16 GB of memory. The proposed IP problems (4.2) and (4.3) are solved through Gurobi 11.0 [46]. From preliminary results, the impact of validating the weight parameter resulted in negligible effects. Thus, in the following experimental evaluation we fixed , ensuring a balanced tradeoff between loss and stability of extracted rules. In order to compare multiple methods in terms of predictive power over the selected datasets, we employed the non-parametric Friedman test [32] combined with the signed-rank Wilcoxon-Holm post-hoc test [49], with a level of statistical significance equal to . According to Benavoli et al. [10], the Wilcoxon signed-rank test is more appropriate than the more common post-hoc Nemenyi test [68]. Details about benchmark datasets are contained in Table 2, Table 5 and Table 8. During data preprocessing, we applied one-hot encoding to tabular data with categorical features and discarded samples containing missing values. For regression tasks, we standardized the response vector. No data preprocessing was applied to time series datasets.
@@ 5.1 Tabular regression
In this experiment, we compare the proposed method to the state-of-the-art about rule extraction for regression tasks of tabular data. In particular, competing methods are RuleFit [40], SIRUS [8] and NodeHarvest (NH) [64]. The experimental plan involves running each method on a selected benchmark of 22 datasets from the OpenML public repository [82]. Following the experimental plan of earlier works [60, 64], we extract rules from a regression forest of 500 trees of depth 3 implemented through the scikit-learn package with default settings [67], and fix the maximum amount of selected rules to 15. Since this quantity coincides with parameter , the proposed validation procedure is unnecessary. Instead, we discard rules that cover fewer than a minimum percentage of the training samples, where is tuned through cross-validation. For competing methods, we fix the number of extracted rules (or nodes) to 15 and the maximal interaction order (or length of rules) to 3. Consequently, NodeHarvest does not need further parameters tuning [64]. For SIRUS, we tune the frequency threshold through cross-validation, and we set the parameter to 10, following the authors’ guidelines [8]. For RuleFit, we tune the regularization parameter for the LASSO regression through cross-validation [40]. We divide each dataset into a 75/25 train-test split. We run each regression task with 30 different random seeds, and we report average and standard deviation of the out-of-sample Mean Squared Error (MSE) in Table 2. From obtained results, we find that differences in average predictive performance are not statistically significant (). We show the pairwise statistical comparison through the critical difference diagram comparing ranks in Figure 5. Remarkably, our approach achieves statistically equivalent performance to RuleFit and SIRUS, while presenting a significant improvement over NodeHarvest in terms of MSE (). The median value of computing times of the proposed method is minutes, and the average runtime of the Gurobi IP solver is of the total computing time. On the selected benchmark, competing methods are more efficient than ours.
@@
Dataset | Our method | RuleFit | SIRUS | NH |
Mercedes_Benz | 0.44 0.10 | 0.46 0.11 | 0.51 0.10 | 0.45 0.10 |
Moneyball | 0.26 0.07 | 0.22 0.08 | 0.20 0.07 | 0.26 0.08 |
abalone | 0.61 0.06 | 0.61 0.06 | 0.68 0.06 | 0.60 0.06 |
autoMpg | 0.31 0.06 | 0.30 0.08 | 0.28 0.05 | 0.32 0.08 |
bank32nh | 0.65 0.03 | 0.64 0.04 | 0.61 0.04 | 0.65 0.04 |
boston | 0.28 0.04 | 0.27 0.06 | 0.29 0.09 | 0.31 0.06 |
cpu_small | 0.08 0.01 | 0.08 0.01 | 0.11 0.01 | 0.08 0.01 |
elevators | 0.52 0.03 | 0.54 0.05 | 0.64 0.04 | 0.53 0.03 |
house_16H | 0.68 0.05 | 0.69 0.06 | 0.66 0.06 | 0.69 0.05 |
houses | 0.51 0.01 | 0.52 0.03 | 0.56 0.02 | 0.51 0.01 |
kin8nm | 0.67 0.02 | 0.68 0.02 | 0.65 0.03 | 0.68 0.05 |
mtp | 0.74 0.03 | 0.73 0.04 | 0.74 0.03 | 0.75 0.06 |
no2 | 0.59 0.09 | 0.57 0.12 | 0.54 0.10 | 0.60 0.13 |
pol | 0.29 0.02 | 0.28 0.05 | 0.28 0.02 | 0.30 0.01 |
puma32H | 0.41 0.02 | 0.42 0.08 | 0.56 0.06 | 0.42 0.01 |
socmob | 0.33 0.09 | 0.32 0.11 | 0.37 0.12 | 0.37 0.25 |
space_ga | 0.53 0.08 | 0.54 0.10 | 0.53 0.10 | 0.56 0.09 |
stock | 0.09 0.01 | 0.12 0.03 | 0.24 0.08 | 0.10 0.01 |
tecator | 0.05 0.01 | 0.06 0.02 | 0.08 0.02 | 0.04 0.02 |
us_crime | 0.46 0.07 | 0.33 0.04 | 0.35 0.09 | 0.50 0.09 |
wind | 0.36 0.01 | 0.36 0.03 | 0.39 0.02 | 0.36 0.03 |
wine_quality | 0.74 0.04 | 0.74 0.04 | 0.78 0.04 | 0.75 0.02 |
Average | 0.44 0.04 | 0.43 0.06 | 0.46 0.06 | 0.45 0.05 |
Rank () | 2.11 | 2.23 | 2.73 | 2.93 |
5.1.1 Comparison in terms of fidelity
To assess internal fidelity, we compute the two representativeness measures introduced in Section 3.4, and the F1-score between the features selected by each rule extraction methods and the top of features ranked by their importance in the random forest, following the work of Velmurugan et al. [84]. To evaluate external fidelity, we report the disagreement in terms of MSE computed between the predictions of each competing method and those of the random forest. We report in Table 3 the average of each fidelity measure computed over the benchmark datasets, for each rule extraction method. For the sake of completion, we also report the average number of extracted rules. Notably, the proposed method obtains the best scores of representativeness and disagreement, while obtaining the second best F1-score. Instead, RuleFit achieves the lowest disagreement value and the best F1-score, and obtains the second best values in terms of representativeness. On average, the proposed method and NodeHarvest extract the lowest amount of rules from the tree ensemble. By contrast, SIRUS extracts the largest number of rules, due to its tendency to select rules with a very limited amount of logical conditions. We conclude that the choice of the interpretable method depends on the desired result. Indeed, SIRUS is the preferred method to obtain a list containing a large amount of short rules. On the contrary, the proposed approach is the most suitable to extract a list that contains a smaller amount of more representative rules. We show an example of extracted rules in B.
Measure | Our method | RuleFit | SIRUS | NH |
Represented trees | 0.731 | 0.607 | 0.307 | 0.514 |
Represented paths | 0.069 | 0.032 | 0.005 | 0.042 |
F1-score | 0.543 | 0.586 | 0.447 | 0.505 |
Disagreement | 0.047 | 0.047 | 0.092 | 0.065 |
Extracted rules | 8.44 | 10.21 | 11.50 | 7.99 |